OpenStudy (anonymous):

How do I write a systems of inequalities? Jason id buying food for a party. One package of chicken wings costs $6.50, and Hot Dogs cost $4.00 per pound. He must spend at least $80.00 (I have 6.5x+4y<80) **Hard part is-He must buy at least 5 pounds of hot dogs, and at least 3 packages of chicken wings.**

OpenStudy (jdoe0001):

@yasmin12380 seems to me that $$ c = chicken\ bags\\ h = hotdogs\\ c \ge 3 \ \ and \ \ h \ge 5\\ 6.5(3c)+4(5h) \ge 80 $$